# High-low analysis and small cap decoupling over non-Archimedean local fields

### v1

@@ 1.2 Initial notation-setting
Let $\mathbb{K}$ be a non-Archimedean local field of characteristic not $2$, i.e. a nondiscrete totally disconnected locally compact topological field which is equipped with a complete absolute value ||$\cdot$|| ⋅ | inducing the topology, for which $2\neq 0$. We normalize ||$\cdot$|| ⋅ | by insisting that it is the modular function for $\mathbb{K}$, regarded as a LCA group. Let $\mathcal{V}\subseteq(0,\infty)$ be the range of ||$\cdot$|| ⋅ | on the nonzero members of $\mathbb{K}$. Let $\mathbf{m}\subseteq\mathcal{O}$ be the maximal ideal, and $\varpi\in\mathbf{m}$ be a uniformizer. Let $\eta\in\mathbb{N}$ be minimal such that $\mathbf{p}=|\varpi|^{-\eta}$ satisfies $\mathbf{p}\geq|2|^{-1}$. Fix some $R\in\mathbf{p}^{2\mathbb{N}}$, and write $N=\frac{1}{2}\log_{\mathbf{p}}(R)$. For $0\leq k\leq N$, we write $R_{k}=\mathbf{p}^{k}$. Let $\alpha\in(0,R)$ and $U_{\alpha}=\{x\in B_{R}:|f(x)|>\alpha\}$.
@@
$M_{a,\lambda}=\begin{bmatrix}1&0\\ a&1\end{bmatrix}.\mathrm{diag}(\lambda,\lambda^{2}).$
@@
$e\left(\sum_{n=N}^{\infty}a_{n}p^{n}\right)=\exp(2\pi ia_{-1}p^{-1}),\quad N\in\mathbb{Z},\{a_{n}\}_{n}\in\{0,\ldots,p-1\}^{\mathbb{Z}}$
@@
[…]
for any Schwartz-Bruhat function $f:\mathbb{K}\to\mathbb{C}$. Functions on $\mathbb{K}^{n}$ will be handled similarly.
@@
$N_{\tau_{I}}=M_{a,1}.\mathrm{diag}(\varpi^{-\eta(N-k)},\varpi^{-\eta N}),$
where we make arbitrary choices of $a\in I$. Let $\mathcal{U}_{\tau_{I}}$ be the image of the set of translates of $\mathcal{O}^{2}$ under $N_{\tau_{I}}$; $\mathcal{U}$ does not depend on the choice of $a$. For $U\in\mathcal{U}_{\tau_{I}}$, define the averaging operator $\mathcal{A}_{U}$ by
[…]
@@ 1.3 Overview of remainder of paper
We begin by establishing a number of technical high/low decomposition results, applicable to a general Schwartz-Bruhat function $f$ with Fourier support in $\mathcal{N}_{R^{-1}}(\mathbb{P}^{1})$. Later, we will fix a single $f$ and perform a pruning procedure, in order to obtain appropriate functions for which the preceding results are useful.
@@ 2.1 Technical lemmas
$\int\Big{|}\sum_{\theta}|f_{\theta}|^{2}*\chi_{>R_{k}/R}^{\vee}\Big{|}^{2}=\int_{|\xi|>R_{k}/R}\Big{|}\sum_{\tau_{k}}\sum_{\theta\prec\tau_{k}}\widehat{|f_{\theta}|^{2}}\Big{|}^{2}.$
@@
$\begin{split}\int\Big{|}\sum_{\tau_{k}}|f_{\tau_{k}}|^{2}*\chi_{>R_{k+l}^{-1}}^{\vee}\Big{|}^{2}&=\int_{\mathbb{Q}_{p}^{2}\setminus B_{R_{k+l}^{-1}}}\Big{|}\sum_{\tau_{k}}\widehat{|f_{\tau_{k}}|^{2}}\Big{|}^{2}\\ &\leq|2|^{-1}\mathbf{p}^{l}\sum_{\tau_{k}}\int\Big{|}|f_{\tau_{k}}|^{2}*\chi_{>R_{k+l}^{-1}}^{\vee}\Big{|}^{2}\end{split}$
@@
$\int\left|\sum_{\tau_{\ell}}|f_{\tau_{\ell}}|^{2}*\chi_{>R_{\ell}/R}^{\vee}\right|^{2}\leq 3\mathbf{p}^{4}N\sum_{k=\ell}^{N}\sum_{\tau_{k}}\int\Big{|}\sum_{\theta\prec\tau_{k}}|f_{\theta}|^{2}*\chi_{\leq R_{k}/R}\Big{|}^{2}.$
@@
Consider the first summand. By the low lemma (2.1),
@@
By the high lemma (a) and Plancherel, we have
@@
Next, for each $\ell+1\leq k\leq N$, by the low lemma (2.1),
@@
$\begin{split}\int\Big{|}\sum_{\tau_{k-1}\prec\tau_{s}}|f_{\tau_{k-1}}|^{2}*\chi_{(R_{k}^{-1},R_{k-1}^{-1}]}^{\vee}\Big{|}^{2}&\leq 2\mathbf{p}\sum_{\tau_{k-1}}\int\Big{|}\sum_{\theta\prec\tau_{k-1}}|f_{\theta}|^{2}\Big{|}^{2}\\ &=2\mathbf{p}\sum_{\tau_{k-1}}\int\Big{|}\sum_{\theta\prec\tau_{k-1}}|f_{\theta}|^{2}*\chi_{\leq R_{k-1}/R}^{\vee}\Big{|}^{2}\\ &+2\mathbf{p}\sum_{k\leq t\leq N}\sum_{\tau_{k-1}}\int\Big{|}\sum_{\theta\prec\tau_{k-1}}|f_{\theta}|^{2}*\chi_{(R_{t-1}/R,R_{t}/R]}^{\vee}\Big{|}^{2}.\end{split}$
@@
By Cauchy-Schwarz and Plancherel, we have
$\sum_{\tau_{t-1}}\int\Big{|}\sum_{\theta\prec\tau_{t-1}}|f_{\theta}|^{2}*\chi_{(R_{t-1}/R,R_{t}/R]}^{\vee}\Big{|}^{2}\leq\mathbf{p}^{3}\sum_{\tau_{t}}\int\Big{|}\sum_{\theta\prec\tau_{t}}|f_{\theta}|^{2}*\chi_{\leq R_{t}/R}^{\vee}\Big{|}^{2}.$
Summing the above bounds, we achieve the desired result.
We conclude this subsection by recording the Córdoba-Fefferman argument over $\mathbb{K}$, which is here represented by two results: Theorem (2.4) (the local form) and Prop. (2.5) (the global form).
@@
Suppose $|2|=\epsilon\in(0,1]$ in $\mathbb{K}$. Let $S\leq D\leq\epsilon\Gamma\leq\epsilon$, and $R>0$. Let $J\in\mathcal{P}(\mathcal{O},\Gamma)$ and $L,L^{\prime}\in\mathcal{P}(J,\epsilon^{-1}D)$ be distinct, $I,I^{\prime}\in\mathcal{P}(L,D),\mathcal{P}(L^{\prime},D)$, respectively, and write $\tau=\tau_{I},\tau^{\prime}=\tau_{I^{\prime}}$ for the caps above $I,I^{\prime}$ on $\mathbb{P}^{1}$. Then we have the estimate
$\int_{B_{R}}|f_{\tau}|^{2}|f_{\tau^{\prime}}|^{2}\leq\max(1,(\Gamma/D)^{2}(SR)^{-2})\int_{B_{R}}\left(\sum_{K\in\mathcal{P}(J,S)}|f_{\tau_{K}}|^{2}\right)^{2}.$
@@
where a given tuple $(\theta_{1},\theta_{2},\vartheta_{1},\vartheta_{2})$ produces a nontrivial summand only if $(\theta_{1}-\theta_{2}+\vartheta_{1}-\vartheta_{2})\cap B_{R^{-1}}\neq\emptyset$. Suppose this is the case. Let $\xi_{1},\xi_{2},\eta_{1},\eta_{2}$ be such that
$(\xi_{1},\xi_{1}^{2})\in\theta_{1},(\xi_{2},\xi_{2}^{2})\in\theta_{2},\quad(\eta_{1},\eta_{1}^{2})\in\vartheta_{1},(\eta_{2},\eta_{2}^{2})\in\vartheta_{2}.$
@@
indeed, the two bracketed expressions belong to distinct members of $\mathcal{P}(\mathcal{O},|2|\epsilon^{-1}D)$. By the assumption on the caps, $|(\xi_{1}-\xi_{2})+(\eta_{1}-\eta_{2})|\leq R^{-1}$. Thus,
@@
The first summand has size $>D|\xi_{1}-\xi_{2}|$, so the vanishing condition may hold only if $|\xi_{1}-\xi_{2}|<D^{-1}\Gamma R^{-1}$. Symmetrically, it is necessary for $|\eta_{1}-\eta_{2}|<D^{-1}\Gamma R^{-1}$.
If $D^{-1}\Gamma R^{-1}\leq S$, we are done. Assume now $D^{-1}\Gamma R^{-1}\geq S$. For each $\theta_{1}\prec\tau$ and $\vartheta_{1}\prec\tau^{\prime}$ we may find $\leq(D^{-1}\Gamma S^{-1}R^{-1})^{2}$ pairs $\theta_{2}\prec\tau,\vartheta_{2}\prec\tau^{\prime}$ for which the corresponding integral is nonzero. Thus, by the Schur test,
$\int_{B_{R}}|f_{\tau}|^{2}|f_{\tau^{\prime}}|^{2}\leq(\Gamma/D)^{2}(SR)^{-2}\int_{B_{R}}\sum_{\theta\prec\tau,\vartheta\prec\tau^{\prime}}|f_{\theta}|^{2}|f_{\vartheta}|^{2}.$
The result follows immediately.
@@
Proposition 2.5 (Córdoba-Fefferman square function estimate).
Suppose $\mathrm{char}(\mathbb{K})\neq 2$. If $f$ has Fourier support in $\mathcal{N}_{\delta}(\mathbb{P}^{1})\subseteq\mathbb{Q}_{p}^{2}$, then
@@ 2.2 Bounding the broad sets
$\mathcal{G}_{\theta}=\left\{U\in\mathcal{U}_{\theta}:\fint_{U}|f_{\theta}|^{2}\geq\frac{\alpha^{2}}{2e^{2}(\#\theta)^{2}}\right\},$
@@
$\mathcal{G}_{\tau_{k}}:=\left\{U\in\mathcal{U}_{\tau_{k}}:\fint_{U}\sum_{\theta\prec\tau_{k}}|f_{k+1,\theta}|^{2}\geq\frac{\alpha^{2}}{2e^{2}(\#\tau_{k})^{2}}\right\},$
@@
$\int\Big{|}\mathcal{A}_{\mathcal{U}_{\tau_{k}}}\Big{[}\sum_{\theta\prec\tau_{k}}|f_{m,\theta}^{\mathcal{B}}|^{2}\Big{]}|\Big{|}^{2}\leq\int\left|\mathcal{A}_{\mathcal{U}_{\tau_{k}}}[g_{k,\tau_{k}}]\right|^{2}.$
@@
$V_{\alpha}=\Big{\{}x\in B_{R}:|f_{N}(x)|>(1-\frac{1}{2^{1/2}eN})\alpha\Big{\}}.$
@@
$\begin{split}V_{\alpha}\subseteq\Big{\{}x\in V_{\alpha}:|f_{0}(x)|\geq(N^{5}+1)^{-1}|&f_{N}(x)|\Big{\}}\cup\bigcup_{m=1}^{N}\Big{\{}x\in V_{\alpha}:|f_{m}^{\mathcal{B}}|(x)\geq N^{-1}(1+N^{-5})^{-1}|f_{N}(x)|\Big{\}}\\ &=:U_{\alpha}^{0}\cup\bigcup_{m=1}^{N}U_{\alpha}^{m}.\end{split}$ (2.2)
@@
$\fint_{B_{R}}\sum_{\theta}|f_{\theta}|^{2}\geq\alpha^{2}.$
@@
$\alpha^{2}\int_{U_{\alpha}^{0}}|f_{0}|^{2}\leq\mu(U)\left(\fint_{B_{R}}\sum_{\theta}|f_{\theta}|^{2}\right)^{2}.$
Finally, on $U_{\alpha}^{0}$ we have $|f_{0}(x)|\geq(1-\frac{1}{2^{1/2}eN})\frac{\alpha}{N^{5}+1}$; the result follows. ∎
@@
$\fint_{U}\sum_{\theta\prec\tau_{\ell}}|f_{\ell,\theta}^{\mathcal{B}}|^{2}*\chi_{\leq R_{\ell}/R}^{\vee}\leq\frac{\alpha^{2}}{2e^{2}(\#\tau_{\ell})^{2}}.$
@@
(b): The assumption and Cauchy-Schwarz give
@@
$\frac{C^{2}\alpha^{2}}{2e^{2}N^{2}}<\frac{C^{2}\alpha^{2}(\#\tau_{\ell}\prec\tau_{k})^{2}}{2e^{2}(\#\tau_{\ell})^{2}},$
@@
Let $1\leq k\leq m\leq N$. Suppose $\tau_{k},\tau_{k}^{\prime}\prec\tau_{k-1}$ are distinct. Write $\ell=\max(m-1,k)$. Then we have
$\int_{\mathrm{Br}_{\alpha}^{m}(\tau_{k},\tau_{k}^{\prime})}|f_{m,\tau_{k}}^{\mathcal{B}}f_{m,\tau_{k}^{\prime}}^{\mathcal{B}}|^{2}\leq 4\mathbf{p}^{2}\int_{\mathbb{Q}_{p}^{2}}\Big{|}\sum_{\tau_{\ell}\prec\tau_{k-1}}|f_{m,\tau_{\ell}}^{\mathcal{B}}|^{2}*\chi_{>R_{\ell}/R}^{\vee}\Big{|}^{2}.$
@@
By bilinear restriction, using the fact that $\mathbf{p}\geq|2|^{-1}$, in either case $k\leq m-1$ or $k\geq m$, we achieve the bound
$\int_{\operatorname{Br}_{\alpha}^{m}(\tau_{k},\tau_{k}^{\prime})}|f_{m,\tau_{k}}^{\mathcal{B}}f_{m,\tau_{k}^{\prime}}^{\mathcal{B}}|^{2}\leq\mathbf{p}^{2}\int_{\operatorname{Br}_{\alpha}^{m}(\tau_{k},\tau_{k}^{\prime})+B(0,R_{\ell})}\Big{|}\sum_{\tau_{\ell}\prec\tau_{k-1}}|f_{m,\tau_{\ell}}^{\mathcal{B}}|^{2}\Big{|}^{2}.$
For each $x\in\mathrm{Br}_{\alpha}^{m}(\tau_{k},\tau_{k}^{\prime})$, we have that $|f_{m,\tau_{k-1}}^{\mathcal{B}}(x)|\geq 2^{1/2}e^{-1}N^{-1}\alpha$. Thus, by the weak high-domination lemma (2.9),
$\sum_{\tau_{\ell}\prec\tau_{k-1}}|f_{m,\tau_{\ell}}^{\mathcal{B}}|^{2}(x)\leq 2\Big{|}\sum_{\tau_{\ell}\prec\tau_{k-1}}|f_{m,\tau_{\ell}}^{\mathcal{B}}|^{2}*\chi_{>R_{\ell}/R}^{\vee}(x)\Big{|}.$

### v2

@@ 1.3 Initial notation-setting
Let $\mathbb{K}$ be a non-Archimedean local field of characteristic not $2$, i.e. a nondiscrete totally disconnected locally compact topological field which is equipped with a complete absolute value $|\cdot|$ inducing the topology, for which $2\neq 0$. We normalize $|\cdot|$ by insisting that it is the modular function for $\mathbb{K}$, regarded as a LCA group.
Let $\mathcal{V}\subseteq(0,\infty)$ be the range of $|\cdot|$ on the nonzero members of $\mathbb{K}$. Let $\mathbf{m}\subseteq\mathcal{O}$ be the maximal ideal, and $\varpi\in\mathbf{m}$ be a uniformizer. Let $\eta\in\mathbb{N}$ be minimal such that $\mathbf{p}=|\varpi|^{-\eta}$ satisfies $\mathbf{p}\geq|2|^{-1}$; note that this $\mathbf{p}$ matches with the choice in (1.2). We will also write $\mathbf{q}=\#\mathbf{k}=p^{f}$ for the cardinality of the residue field $\mathbf{k}$ of $\mathbb{K}$. Note in particular that $\mathbf{q}=\mathbf{p}$ if $\mathrm{char}(\mathbf{k})\neq 2$, and $\mathbf{q}\leq\mathbf{p}=|2|^{-1}$ if $\mathrm{char}(\mathbf{k})=2$.
Fix some $R\in\mathbf{p}^{2\mathbb{N}}$, and write $N=\frac{1}{2}\log_{\mathbf{p}}(R)$. For $0\leq k\leq N$, we write $R_{k}=\mathbf{p}^{k}$. Let $\alpha\in(0,R)$ and $U_{\alpha}=\{x\in B_{R}:|f(x)|>\alpha\}$.
@@
$M_{a,\lambda}=\begin{bmatrix}1&0\\ 2a&1\end{bmatrix}.\mathrm{diag}(\lambda,\lambda^{2}).$
@@
$e\left(\sum_{n=N}^{\infty}a_{n}p^{n}\right)=\exp\left(2\pi i\sum_{n=N}^{-1}a_{n}p^{n}\right),\quad N\in\mathbb{Z},\{a_{n}\}_{n}\in\{0,\ldots,p-1\}^{\mathbb{Z}}$ (1.10)
@@
$\hat{f}(\xi)=\int_{\mathbb{K}}e(x\xi)f(x)d\mu(x),$
for any Schwartz–Bruhat function $f:\mathbb{K}\to\mathbb{C}$. Functions on $\mathbb{K}^{n}$ will be handled similarly.
@@
$N_{\tau_{I}}=M_{a,1}^{-\top}.\mathrm{diag}(\varpi^{-\eta(N-k)},\varpi^{-\eta N}),$
where we make arbitrary choices of $a\in I$. Let $\mathcal{U}_{\tau_{I}}$ be the image of the set of translates of $\mathcal{O}^{2}$ under $N_{\tau_{I}}$; $\mathcal{U}_{\tau_{I}}$ does not depend on the choice of $a$. For $U\in\mathcal{U}_{\tau_{I}}$, define the averaging operator $\mathcal{A}_{U}$ by
$\mathcal{A}_{U}[f]=\mu(U)^{-1}\int_{U}fd\mu.$
@@ 1.4 Overview of remainder of paper
We begin by establishing a number of technical high/low decomposition results, applicable to a general Schwartz–Bruhat function $f$ with Fourier support in $\mathcal{N}_{R^{-1}}(\mathbb{P}^{1})$. Later, we will fix a single $f$ and perform a pruning procedure, in order to obtain appropriate functions for which the preceding results are useful.
@@ 2.1 Technical lemmas
$\int\Big|\sum_{\theta}|f_{\theta}|^{2}*\chi_{>R_{k}/R}^{\vee}\Big|^{2}=\int_{|\xi|>R_{k}/R}\Big|\sum_{\tau_{k}}\sum_{\theta\prec\tau_{k}}\widehat{|f_{\theta}|^{2}}(\xi)\Big|^{2}.$
@@
$\begin{split}\int\Big|\sum_{\tau_{k}}|f_{\tau_{k}}|^{2}*\chi_{>R_{k+l}^{-1}}^{\vee}\Big|^{2}&=\int_{\mathbb{Q}_{p}^{2}\setminus B_{R_{k+l}^{-1}}}\Big|\sum_{\tau_{k}}\widehat{|f_{\tau_{k}}|^{2}}\Big|^{2}\\ &\leq|2|^{-1}\mathbf{p}^{l}p^{-f}\sum_{\tau_{k}}\int\Big||f_{\tau_{k}}|^{2}*\chi_{>R_{k+l}^{-1}}^{\vee}\Big|^{2}\end{split}$
@@
$\int\left|\sum_{\tau_{\ell}}|f_{\tau_{\ell}}|^{2}*\chi_{>R_{\ell}/R}^{\vee}\right|^{2}\leq 3\mathbf{q}^{-1}\mathbf{p}^{3}N\sum_{k=\ell+1}^{N}\sum_{\tau_{k}}\int\Big|\sum_{\theta\prec\tau_{k}}|f_{\theta}|^{2}*\chi_{\leq R_{k}/R}\Big|^{2}.$
@@
Consider the first sum. By the low lemma (2.1),
@@
By the high lemma (a) and Cauchy–Schwarz, we have
@@
Next, we consider the second sum in our decomposition. For each $\ell+1\leq k\leq N$, by the low lemma (2.1),
@@
$\begin{split}\int\Big|\sum_{\tau_{k-1}}|f_{\tau_{k-1}}|^{2}*\chi_{(R_{k}^{-1},R_{k-1}^{-1}]}^{\vee}\Big|^{2}&\leq 2|2|^{-1}\mathbf{q}^{-1}\mathbf{p}\sum_{\tau_{k-1}}\int\Big|\sum_{\theta\prec\tau_{k-1}}|f_{\theta}|^{2}\Big|^{2}\\ &=2|2|^{-1}\mathbf{q}^{-1}\mathbf{p}\sum_{\tau_{k-1}}\int\Big|\sum_{\theta\prec\tau_{k-1}}|f_{\theta}|^{2}*\chi_{\leq R_{k-1}/R}^{\vee}\Big|^{2}\\ &+2|2|^{-1}\mathbf{q}^{-1}\mathbf{p}\sum_{k\leq t\leq N}\sum_{\tau_{k-1}}\int\Big|\sum_{\theta\prec\tau_{k-1}}|f_{\theta}|^{2}*\chi_{(R_{t-1}/R,R_{t}/R]}^{\vee}\Big|^{2}.\end{split}$
@@
By Cauchy–Schwarz, we have
$\sum_{\tau_{t-1}}\int\Big|\sum_{\theta\prec\tau_{t-1}}|f_{\theta}|^{2}*\chi_{(R_{t-1}/R,R_{t}/R]}^{\vee}\Big|^{2}\leq\mathbf{p}\sum_{\tau_{t}}\int\Big|\sum_{\theta\prec\tau_{t}}|f_{\theta}|^{2}*\chi_{\leq R_{t}/R}^{\vee}\Big|^{2}.$
Adding our two contributions, we obtain
$\int\left|\sum_{\tau_{\ell}}|f_{\tau_{\ell}}|^{2}*\chi_{>R_{\ell}/R}^{\vee}\right|^{2}\leq\sum_{k=\ell+1}^{N}\Big(\frac{2\mathbf{p}^{2}}{|2|\mathbf{q}}(k-\ell)+\mathbf{p}\Big)\sum_{\tau_{k}}\int\Big|\sum_{\theta\prec\tau_{k}}|f_{\theta}|^{2}*\chi_{\leq R_{k}/R}\Big|^{2}.$
We conclude this subsection by recording the Córdoba–Fefferman argument over $\mathbb{K}$, which is here represented by two results: Theorem (2.4) (the local form) and Prop. (2.5) (the global form).
@@
Suppose $|2|=\epsilon\in(0,1]$ in $\mathbb{K}$. Let
$0<S\leq D\leq\epsilon\Gamma\leq\epsilon,\quad R>0.$
Let $J\in\mathcal{P}(\mathcal{O},\Gamma)$ and $L,L^{\prime}\in\mathcal{P}(J,\epsilon^{-1}D)$ be distinct, $I,I^{\prime}\in\mathcal{P}(L,D),\mathcal{P}(L^{\prime},D)$, respectively, and write $\tau=\tau_{I},\tau^{\prime}=\tau_{I^{\prime}}$ for the caps above $I,I^{\prime}$ on $\mathbb{P}^{1}$. Then we have the estimate
$\int_{B_{R}}|f_{\tau}|^{2}|f_{\tau^{\prime}}|^{2}\leq\max(\mathbf{q}^{-2}(\Gamma/D)^{2}(SR)^{-2},1)\int_{B_{R}}\left(\sum_{K\in\mathcal{P}(J,S)}|f_{\tau_{K}}|^{2}\right)^{2}.$
@@
where a given tuple $(\theta_{1},\theta_{2},\vartheta_{1},\vartheta_{2})$ produces a nontrivial summand only if $(\theta_{1}-\theta_{2}+\vartheta_{1}-\vartheta_{2})\cap B_{R^{-1}}\neq\emptyset$. Suppose this is the case. We may find $\xi_{1},\xi_{2},\eta_{1},\eta_{2}$ be such that
$(\xi_{1},\xi_{1}^{2})\in\theta_{1},(\xi_{2},\xi_{2}^{2})\in\theta_{2},\quad(\eta_{1},\eta_{1}^{2})\in\vartheta_{1},(\eta_{2},\eta_{2}^{2})\in\vartheta_{2},$
and such that there are $\varepsilon_{1},\varepsilon_{2},\delta_{1},\delta_{2}\in\mathbb{K}$ with $|\varepsilon_{\iota}|,|\delta_{\iota}|\leq S^{2}$, such that
$(\xi_{1},\xi_{1}^{2}+\varepsilon_{1})-(\xi_{2},\xi_{2}^{2}+\varepsilon_{2})+(\eta_{1},\eta_{1}^{2}+\delta_{1})-(\eta_{2},\eta_{2}^{2}+\delta_{2})\in B_{R^{-1}}.$
@@
indeed, the two bracketed expressions belong to distinct members of $\mathcal{P}(\mathcal{O},|2|\epsilon^{-1}D)$. By the assumption on the points, $|(\xi_{1}-\xi_{2})+(\eta_{1}-\eta_{2})|\leq R^{-1}$. Thus,
@@
The first summand has size $D|\xi_{1}-\xi_{2}|$, so the vanishing condition may hold only if
$|\xi_{1}-\xi_{2}|<D^{-1}\max(\Gamma R^{-1},S^{2}).$
Symmetrically, it is necessary for $|\eta_{1}-\eta_{2}|<T:=D^{-1}\max(\Gamma R^{-1},S^{2})$. We divide into cases. Suppose first that $T\leq\mathbf{q}S$. Then we certainly have
$\int_{B_{R}}|f_{\tau}|^{2}|f_{\tau^{\prime}}|^{2}=\sum_{K_{1}\in\mathcal{P}(I,S)}\sum_{K_{2}\in\mathcal{P}(I^{\prime},S)}\int_{B_{R}}|f_{\tau_{K_{1}}}|^{2}|f_{\tau_{K_{2}}}|^{2},$
and we are done.
In the alternate case, we have $S<\mathbf{q}^{-1}T$, and we take our weaker identity
$\int_{B_{R}}|f_{\tau}|^{2}|f_{\tau^{\prime}}|^{2}=\sum_{J_{1}\in\mathcal{P}(I,\mathbf{q}^{-1}T)}\sum_{J_{2}\in\mathcal{P}(I^{\prime},\mathbf{q}^{-1}T)}\int_{B_{R}}|f_{\tau_{J_{1}}}|^{2}|f_{\tau_{J_{2}}}|^{2},$
and apply Cauchy–Schwarz to conclude
$\int_{B_{R}}|f_{\tau}|^{2}|f_{\tau^{\prime}}|^{2}\leq\mathbf{q}^{-2}S^{-2}T^{2}\sum_{K_{1}\in\mathcal{P}(I,T)}\sum_{K_{2}\in\mathcal{P}(I^{\prime},T)}\int_{B_{R}}|f_{\tau_{K_{1}}}|^{2}|f_{\tau_{K_{2}}}|^{2}.$
Finally, we have the implication
$S<\mathbf{q}^{-1}T\implies S^{2}\leq\Gamma R^{-1},$
so that our inequalities may take the shape
$\int_{B_{R}}|f_{\tau}|^{2}|f_{\tau^{\prime}}|^{2}\leq\int_{B_{R}}\left(\sum_{K\in\mathcal{P}(J,S)}|f_{\tau_{K}}|^{2}\right)^{2}\times\begin{cases}1&T\leq\mathbf{q}S\\ \mathbf{q}^{-2}(\Gamma/D)^{2}(RS)^{-2}&S<\mathbf{q}^{-1}T,\end{cases}$
which clearly implies our conclusion.
@@
Proposition 2.5 (Córdoba–Fefferman square function estimate).
Suppose $\mathrm{char}(\mathbb{K})\neq 2$. If $f$ has Fourier support in $\mathcal{N}_{\delta}(\mathbb{P}^{1})\subseteq\mathbb{K}^{2}$, then
@@ 2.2 Bounding the broad sets
$\mathcal{G}_{\theta}=\left\{U\in\mathcal{U}_{\theta}:\fint_{U}|f_{\theta}|^{2}\geq\frac{\alpha^{2}}{2e^{2}N^{2}(\#\theta)^{2}}\right\},$
@@
$\mathcal{G}_{\tau_{k}}:=\left\{U\in\mathcal{U}_{\tau_{k}}:\fint_{U}\sum_{\theta\prec\tau_{k}}|f_{k+1,\theta}|^{2}\geq\frac{\alpha^{2}}{2e^{2}N^{2}(\#\tau_{k})^{2}}\right\},$
@@
Remark 2.7 (Pruning inequalities).
For each $0\leq k<N$ and each $\theta$, we have the pointwise inequalities
$\max\big(|f_{k,\theta}|,|f_{k,\theta}^{\mathcal{B}}|\big)\leq|f_{k+1,\theta}|\leq|f_{\theta}|.$ (2.1)
@@
$\int\Big|\mathcal{A}_{\mathcal{U}_{\tau_{k}}}\Big[\sum_{\theta\prec\tau_{k}}|f_{m,\theta}^{\mathcal{B}}|^{2}\Big]\Big|^{2}\leq\int\left|\mathcal{A}_{\mathcal{U}_{\tau_{k}}}[g_{k,\tau_{k}}]\right|^{2}.$
@@
$V_{\alpha}=\Big\{x\in B_{R}:|f_{N}(x)|>\Big(1-\frac{1}{2^{1/2}eN}\Big)\alpha\Big\}.$
@@
$\begin{split}V_{\alpha}\subseteq\Big\{x\in V_{\alpha}:|f_{0}(x)|\geq(N^{4}+1)^{-1}|&f_{N}(x)|\Big\}\cup\bigcup_{m=1}^{N}\Big\{x\in V_{\alpha}:|f_{m}^{\mathcal{B}}|(x)\geq N^{-1}(1+N^{-4})^{-1}|f_{N}(x)|\Big\}\\ &=:U_{\alpha}^{0}\cup\bigcup_{m=1}^{N}U_{\alpha}^{m}.\end{split}$ (2.3)
@@
$\fint_{B_{R}}\sum_{\theta}|f_{\theta}|^{2}\geq\frac{\alpha^{2}}{2e^{2}N^{2}}.$
@@
$\frac{\alpha^{2}}{2e^{2}N^{2}}\int_{U_{\alpha}^{0}}|f_{0}|^{2}\leq\mu(U)\left(\fint_{B_{R}}\sum_{\theta}|f_{\theta}|^{2}\right)^{2}.$
Finally, on $U_{\alpha}^{0}$ we have $|f_{0}(x)|\geq(1-\frac{1}{2^{1/2}eN})\frac{\alpha}{N^{4}+1}$; the result follows. ∎
@@
$\fint_{U}\sum_{\theta\prec\tau_{\ell}}|f_{\ell,\theta}^{\mathcal{B}}|^{2}*\chi_{\leq R_{\ell}/R}^{\vee}\leq\frac{\alpha^{2}}{2e^{2}N^{2}(\#\tau_{\ell})^{2}}.$
@@
(b): The assumption and Cauchy–Schwarz give
@@
$\frac{C^{2}\alpha^{2}}{2e^{2}N^{2}}<\frac{C^{2}\alpha^{2}N^{2}(\#\tau_{\ell}\prec\tau_{k})^{2}}{2e^{2}(\#\tau_{\ell})^{2}},$
@@
Let $N>2$. Let $1\leq k\leq m\leq N$. Suppose $\tau_{k},\tau_{k}^{\prime}\prec\tau_{k-1}$ are distinct. Write $\ell=\max(m-1,k)$. Then we have
$\int_{\mathrm{Br}_{\alpha}^{m}(\tau_{k},\tau_{k}^{\prime})}|f_{m,\tau_{k}}^{\mathcal{B}}f_{m,\tau_{k}^{\prime}}^{\mathcal{B}}|^{2}\leq 6\mathbf{q}^{-2}\mathbf{p}^{2}\int_{\mathbb{Q}_{p}^{2}}\Big|\sum_{\tau_{\ell}\prec\tau_{k-1}}|f_{m,\tau_{\ell}}^{\mathcal{B}}|^{2}*\chi_{>R_{\ell}/R}^{\vee}\Big|^{2}.$
@@
By local bilinear restriction (2.4), using the fact that $\mathbf{p}\geq|2|^{-1}$, in either case $k\leq m-1$ or $k\geq m$, we achieve the bound
$\int_{\operatorname{Br}_{\alpha}^{m}(\tau_{k},\tau_{k}^{\prime})}|f_{m,\tau_{k}}^{\mathcal{B}}f_{m,\tau_{k}^{\prime}}^{\mathcal{B}}|^{2}\leq\mathbf{q}^{-2}\mathbf{p}^{2}\int_{\operatorname{Br}_{\alpha}^{m}(\tau_{k},\tau_{k}^{\prime})+B(0,R_{\ell})}\Big|\sum_{\tau_{\ell}\prec\tau_{k-1}}|f_{m,\tau_{\ell}}^{\mathcal{B}}|^{2}\Big|^{2}.$
For each $x\in\mathrm{Br}_{\alpha}^{m}(\tau_{k},\tau_{k}^{\prime})$, we have that $|f_{m,\tau_{k-1}}^{\mathcal{B}}(x)|\geq 0.915\cdot e^{-1}N^{-1}\alpha$. Thus, by the weak high-domination lemma (2.10),
$\sum_{\tau_{\ell}\prec\tau_{k-1}}|f_{m,\tau_{\ell}}^{\mathcal{B}}|^{2}(x)\leq 2.5\Big|\sum_{\tau_{\ell}\prec\tau_{k-1}}|f_{m,\tau_{\ell}}^{\mathcal{B}}|^{2}*\chi_{>R_{\ell}/R}^{\vee}(x)\Big|.$